# Calderon-preconditioned boundary integral equations of the Burton–Miller type for transmission problems

### v1

@@ Abstract
This paper proposes well-conditioned boundary integral equations based on the Burton–Miller method for solving transmission problems. The Burton–Miller method is widely accepted as a highly accurate numerical method based on the boundary integral equation for solving exterior wave problems. While this method can also be applied to solve the transmission problems, a straightforward formulation may yield ill-conditioned integral equations. Consequently, the resulting linear algebraic equations may involve a coefficient matrix with a huge condition number, leading to poor convergence of Krylov-based linear solvers. To address this challenge, our study enhances Burton–Miller-type boundary integral equations tailored for transmission problems by exploiting the Calderon formula. In cases where a single material exists in an unbounded host medium, we demonstrate the formulation of the boundary integral equation such that the underlying integral operator ${\cal A}$ is spectrally well-conditioned. Specifically, ${\cal A}$ can be designed in a simple manner that ensures ${\cal A}^{2}$ is bounded and has only a single eigenvalue accumulation point. Furthermore, we extend our analysis to the multi-material case, proving that the square of the proposed operator has only a few eigenvalues except for a compact perturbation. Through numerical examples of several benchmark problems, we illustrate that our formulation reduces the iteration number required by iterative linear solvers, even in the presence of material junction points—locations where three or more sub-domains meet on the boundary.
Keywords: boundary integral equation, the Burton–Miller method, Calderon’s preconditioning, transmission problem, wave scattering, resonance
@@ 1 Introduction
Given that the transmission problem is defined in an unbounded domain, it may be reasonable to use the boundary integral formulation to solve the problem rather than the solver of the domain type such as the finite element method (FEM). Recall that the boundary element method (BEM) bypasses the need for special considerations on the boundary condition at infinity, automatically satisfying the radiation condition.
@@
We assert that the most widely used integral formulations for transmission problems are the PMCHWT (Poggio–Miller–Chang–Harrington–Wu–Tsai) [12] and Müller [13] formulations, both of which are free from real-valued fictitious eigenfrequency. The former is categorised as the integral equation of the first kind, which is thus often combined with some preconditioning techniques. Calderon’s preconditioning [14, 15] is particularly effective in improving the spectral properties of the PMCHWT equations. Niino and Nishimura pointed out that the preconditioning effect can be harnessed by a mere reordering of integral operators [16]. They also showed that just reordering is more effective in reducing total computational time than multiplying the preconditioner when using GMRES. The Müller formulation yields a second-kind integral equation [13], allowing for rapid convergence using the Krylov method. It is, however, noteworthy that these two integral equations may have complex-valued fictitious eigenvalue with a tiny imaginary part in its magnitude. Another possibility is the so-called single boundary integral equation (SBIE) method [17, 18]. For this type of formulation of the second kind, one may augment a parameter roughly controlling the position of the complex-valued fictitious eigenvalue [19]. The SBIE is thus considered promising in both efficiency and accuracy aspects, but applying this to multi-material cases may need some more elaboration [20]. The so-called multi-trace boundary integral formulation [21] would also be an attractive choice, offering a straightforward extension to the multi-material problem. The formulation, however, requires multiple unknown densities on a single boundary element resulting in slow matrix-vector multiplication in a Krylov iteration. A recently proposed formulation [22] involving both the direct and indirect boundary integrals is also efficient especially when the material constant for inclusion is considerably different from that for exterior host medium. The method may suffer from the fictitious eigenvalue problem even for real-valued frequency, though. The Burton–Miller formulation [6], originally designed for exterior boundary value problems, can also be applied to solve the transmission problem. Although the numerical solution by this formulation with a parameter for controlling the fictitious eigenvalue is accurate, a naive use may end up with highly ill-conditioned algebraic equations. With these discussions, we may conclude that there exists no simple, efficient, accurate, and versatile boundary integral formulation for the transmission problem.
@@
    can easily be applied to multi-material cases even with a material junction,
@@
    and has a good spectral characteristic; the square of the underlying integral operator has only a few eigenvalue accumulation points.
@@
The rest of the paper is organised as follows. In Section 2, we state the boundary value problem (BVP) of interest. We then show a straightforward application of the Burton–Miller method to solve the BVP ends up with an ill-conditioned boundary integral formalism and discuss a design concept for integral formulation with good spectral characteristics in Section 3. Sections 4 and 5, the main contributions of the present paper, introduce novel Burton–Miller-type integral equations for single-material and multi-material cases, respectively. After verifying the performance of the present formulations through various numerical experiments in Section 6, we conclude the paper and discuss future directions in Section 7.
@@ 2 Statement of the boundary value problem of interest
where $k_{i}:=\omega\sqrt{\varepsilon_{i}}$ is the wave number of the waves propagating in $\Omega_{i}$, and $\omega>0$ is the underlying angular frequency. Here, we assume the time dependency of all physical quantities in the time domain as $\Re[f(\boldsymbol{x})e^{-\mathrm{i}\omega t}]$, where $f$ is either $u$ or $w$, $\mathrm{i}$ is the imaginary unit, and $t$ is the time. The boundary conditions (2) and (3) on $\Gamma:=\bigcup_{i}^{M}\overline{\partial\Omega_{i}}$ define the jump relations in the total field and its flux. The flux here is defined as $w:=\frac{1}{\varepsilon_{i}}\nabla u\cdot\boldsymbol{n}$ with the material constants $\varepsilon_{i}$, where $\boldsymbol{n}$ is the unit normal on $\Gamma$. $\boldsymbol{n}$ is directed from $\Omega_{1}$ on $\partial\Omega_{1}$. The direction of $\boldsymbol{n}$ on the other boundaries is properly defined and fixed. This type of boundary value problem is often referred to as the transmission problem and is commonly found in engineering fields such as acoustics and electromagnetics. For example, a solution $u$ of the BVP (1)–(4) physically indicates the time-dependent out-of-plane component of a magnetic field, and $w$ corresponds to the in-plane tangential electric wave on the boundary. In this case, $\varepsilon_{i}$ represents the relative dielectric constant of a material filled in $\Omega_{i}$.
@@ 3 Design concepts for the present boundary integral formalism
Since the BVP at hand is defined in an unbounded domain, the boundary integral method is a natural choice for solving it. There are various candidates for the integral formulations, and the choice may significantly affect the accuracy and efficiency of the practical computation. In this study, we ensure the accuracy by utilising the Burton–Miller method. In this section, we discuss the design concept of the integral formulation from the viewpoint of efficiency. Here, let us abstractly denote the integral equation as ${\cal A}y=b$, where $\cal A$ is an integral operator, $y$ and $b$ are respectively unknown and known density functions. The discretised version of the equation upon the collocation, i.e. the system of algebraic linear equations corresponding to the integral equation is denoted as $\mathsf{A}\boldsymbol{y}=\boldsymbol{b}$ with $\mathsf{A}$, $\boldsymbol{y}$, and $\boldsymbol{b}$ respectively being the coefficient matrix, and unknown and known vectors.
@@ 3.1 Integral equation of the second-kind
The efficiency of the BEM largely depends on the spectral property of $\cal A$. It is thus natural to try constructing the integral equation of the Fredholm second kind, i.e. exploring $\cal A$ such that $\cal A=\cal I+\cal K$ holds, where $\cal I$ and $\cal K$ are the identity and compact operators, respectively. The Müller formulation and the SBIE are categorised into this type. It may, unfortunately, be impossible to find such a well-conditioned $\cal A$ if one employs the Burton–Miller method.
@@ 3.2 Integral equation of the first-kind and its Calderon’s preconditioning
Another possibility may take the help of so-called preconditioning. Among the various preconditioning techniques, the Calderon preconditioning [14, 15] is recognised as one of the best for the BEM. This method converts (or preconditions) the original integral equation as ${\cal A}{\cal M}^{-1}z=b$ such that ${\cal A}{\cal M}^{-1}={\cal I}+{\cal K}$ holds, where $\cal M$ is the right preconditioning operator, and $z:={\cal M}y$ is an auxiliary function. Niino and Nishimura [16] demonstrated that, for transmission problems, one may construct, with the PMCHWT formulation, $\cal A$ and $\cal M$ such that ${\cal M}^{-1}\simeq{\cal A}\Leftrightarrow{\cal A}^{2}={\cal I}+{\cal K}$ holds. They also pointed out that, once the integral operator is appropriately formulated in this manner, one does not have to explicitly multiply the (inverse of) preconditioner. This can be understood as follows: Let us consider the preconditioned system $\mathsf{A}^{2}\boldsymbol{z}=\boldsymbol{b}$ upon discretisation and solve this by the GMRES [9]. Let us first recall the fact that GMRES after $n$ iterations finds the best approximated solution of the algebraic equations from the following Krylov subspace:
@@
In summary, we just need to appropriately arrange the integral operator $\cal A$ such that it satisfies ${\cal A}^{2}=\cal I+\cal K$ and to solve the discretised linear equations without multiplying the preconditioner. The present paper is based on the discovery that we can construct such a $\cal A$ for the Burton–Miller-based integral equation even for transmission problems.
@@ 3.3 Spectral property of the conventional Burton–Miller method for transmission problems
where ${\cal I}$ is the identity operator. The integral operator in (20) is notoriously ill-conditioned. The upper and the lower diagonal parts are dominated by the hypersingular operator ${\cal N}$ and the compact one ${\cal S}$, respectively. The eigenvalues of the former grow in magnitude, while those of the latter accumulate at the origin in the complex plane. This high contrast in the eigenvalues makes the resulting coefficient matrix considerably ill-conditioned. In the next section, we modify the integral equation (20) with the help of the Calderon formulae to have a better conditioned one.
4 The present formulation: A single material case
Before deriving an well-conditioned Burton–Miller equations for transmission problems, let us review the properties of the integral operators (9)–(12). First, the operators ${\cal S}$, ${\cal D}$, and ${\cal D}^{*}$ are all compact operators, i.e., their eigenvalues accumulate only at zero in the complex plane. On the other hand, ${\cal N}$ is hypersingular and its eigenvalue may have a large magnitude. It is also well known that the operators have the following relations:
@@ 4 The present formulation: A single material case
$\frac{\cal I}{4}$ $=({\cal D}^{*}_{i})^{2}-{\cal N}_{i}{\cal S}_{i},$ (24)
for $i=1,\cdots,M$. The relations (21)–(24) are known as the Calderon formulae. It is also known that the formulae are valid modulo a compact operator even when the wave numbers of two operators in a single term are different from each other, i.e. the following relation, for example, holds:
$\frac{\cal I}{4}=-{\cal S}_{i}{\cal N}_{j}+{\cal D}_{i}{\cal D}_{j}+{\cal K},$ (25)
for $j\neq i$, where $\cal K$ is a compact operator.
@@
${\cal A}_{1}^{2}$ $=\begin{pmatrix}\frac{1}{4}{\cal I}+\alpha\varepsilon_{2}{\cal S}_{2}{\cal N}_{1}&0\\ \frac{1+\alpha}{4}{\cal I}+\frac{\alpha(1+\alpha)}{2}{\cal N}_{1}-\alpha({\cal N}_{1}{\cal D}_{2}+\alpha{\cal D}_{1}^{*}{\cal N}_{1})-\alpha{\cal S}_{1}{\cal N}_{1}&\frac{\alpha^{2}}{4}{\cal I}+\alpha\varepsilon_{2}{\cal N}_{1}{\cal S}_{2}\end{pmatrix}+{\cal K}$ (35)
@@
where we abusively used ${\cal K}$ to denote two compact operators different from each other. The last equality of (38) is due to the Calderon formulae (21)–(25). Although ${\cal A}_{1}^{2}$ still involves the hypersingular operator ${\cal N}_{1}$, owing to the block lower structure (except for a compact perturbation), its eigenvalue accumulates only at the following two points in the complex plane: $\frac{1-\alpha\varepsilon_{2}}{4}$ and $\frac{\alpha(\alpha-\varepsilon_{2})}{4}$, and they are away from the origin of the complex plane (Recall that the setting $\alpha=-\mathrm{i}/k_{1}$ is used here). With these observations, we may expect that the GMRES applied to the linear algebraic equations obtained by discretising (32) converges fast. Nonetheless, we cannot deny that the hypersingular operator ${\cal N}_{1}$ in ${\cal A}_{1}^{2}$ may affect the convergence of the linear solver.
To derive a better conditioned integral equation, we multiply a complex constant to the first row of (32) as

### v2

# Automatically well-conditioned collocation boundary element method for transmission problems based on the Burton–Miller formulation

@@ Abstract
This paper proposes a collocation boundary element method based on the Burton–Miller method for solving transmission problems, which is rapidly convergent within the Krylov subspace solver framework. Our study enhances Burton–Miller-type boundary integral equations tailored for transmission problems by exploiting the Calderon formula. In cases where a single material exists in an unbounded host medium, we demonstrate the formulation of the boundary integral equation such that the underlying integral operator ${\cal A}$ is spectrally well-conditioned. Specifically, ${\cal A}$ can be designed such that ${\cal A}^{2}$ has only a single eigenvalue accumulation point. Furthermore, we extend this to the multi-material case, proving that the square of the proposed operator has only a few eigenvalues clustering points. When the collocation method is used to discretise the proposed boundary integral equations, the good spectral properties of the integral operator are naturally inherited to the coefficient matrix $\mathsf{A}$ similarly to the Nystöm methods; Almost all eigenvalues of $\mathsf{A}^{2}$ cluster at a few points in the complex plane ensuring the small condition number for $\mathsf{A}$. Through numerical examples of several benchmark problems, we illustrate that our formulation reduces the iteration number required by iterative linear solvers, even in the presence of material junction points.
Keywords: boundary integral equation, collocation BEM, the Burton–Miller method, Calderon’s preconditioning, transmission problem, wave scattering, resonance
@@ 1 Introduction
Given that the transmission problem is defined in an unbounded domain, it may be reasonable to use the boundary integral formulation to solve the problem rather than the solver of the domain type such as the finite element method (FEM). Recall that the boundary element method (BEM) bypasses the need for special considerations on the boundary condition at infinity since it automatically gives solutions satisfying the radiation condition.
@@
We assert that the most widely used integral formulations for transmission problems are the PMCHWT (Poggio–Miller–Chang–Harrington–Wu–Tsai) [12] and Müller [13] formulations, both of which are free from real-valued fictitious eigenfrequency. The former is categorised as the integral equation of the first kind, which is thus often combined with some preconditioning techniques. Calderon’s preconditioning [14, 15] is particularly effective in improving the spectral properties of the PMCHWT equations. Niino and Nishimura pointed out that the preconditioning effect can be harnessed by a mere reordering of integral operators [16]. They also showed that, upon collocation discretisation, just reordering is more effective in reducing total computational time than multiplying the preconditioner when using GMRES. The Müller formulation yields a second-kind integral equation [13], allowing for rapid convergence for the Krylov method. It is, however, noteworthy that these two integral equations may have complex-valued fictitious eigenvalue with a tiny imaginary part in its magnitude. Another possibility is the so-called single boundary integral equation (SBIE) method [17, 18]. For this type of formulation of the second kind, one may augment a parameter roughly controlling the position of the complex-valued fictitious eigenvalue [19]. The SBIE is thus considered promising in both efficiency and accuracy aspects, but applying this to multi-material cases may need some more elaboration [20]. The so-called multi-trace boundary integral formulation [21] would also be an attractive choice, offering a straightforward extension to the multi-material problem. The formulation, however, requires multiple unknown densities on a single boundary element. Hence it results in slow matrix-vector multiplication in a Krylov iteration. A recently proposed formulation [22] involving both the direct and indirect boundary integrals is also efficient especially when the material constant for inclusion is considerably different from that for exterior host medium. The method may suffer from the fictitious eigenvalue problem even for real-valued frequency, though. The Burton–Miller formulation [6], originally designed for exterior boundary value problems, can also be applied to solve the transmission problem. Although the numerical solution by this formulation with a parameter for controlling the fictitious eigenvalue is accurate, a naive use may end up with highly ill-conditioned algebraic equations. With these discussions, we may conclude that there exists no simple, efficient, accurate, and versatile boundary integral formulation for the transmission problem.
@@
    can easily be applied to multi-material cases even with material junctions,
@@
    and has a good spectral characteristic; the square of the underlying integral operator (implicitly computed in GMRES) has only a few eigenvalue accumulation points.
@@
We also mention that we have two options in discretising the boundary integral equation, i.e. the Galerkin and collocation methods. While the former fits well with theoretical discussions based on functional analysis, it requires computationally heavy numerical integrations in the practical implementation. Furthermore, the finite-dimensional matrix obtained as the Galerkin discretised integral operator does not necessarily inherit the spectral properties as is. We thus need some careful discussions with the dual basis and the Gram matrices when deriving the Calderon-preconditioned BEM; see discussions in e.g. [16, 23, 24]. On the other hand, although it is somewhat inferior to the Galerkin method in terms of accuracy, the collocation BEM does not require any special considerations during preconditioning [16]. This is analogous to the fact that Calderon’s preconditioning can be applied seamlessly, even when using the Nyström method [25]. Furthermore, the collocation BEM is attractive from the perspective of simple implementation and easy use [26]. Many engineers thus employ the collocation in practical computations [27, 28]. This is also encouraged by the fact that the hypersingular integrals in BEM can appropriately be evaluated via well-established regularisation techniques [29]. This paper thus discusses the collocation BEM.
The rest of the paper is organised as follows. In Section 2, we state the boundary value problem (BVP) of interest. We then show a straightforward application of the Burton–Miller method to solve the BVP ends up with an ill-conditioned boundary integral formalism and discuss a design concept for integral formulation with good spectral characteristics in Section 3. Section 4 reviews the well-known Calderon formula for boundary integral operators. We also numerically check that the collocation-discretised version of the operators, i.e. matrices, approximately satisfies the same relation as the Calderon formulae even when the operators are defined on the boundary with corners. Sections 5 and 6, the main contributions of the present paper, introduce novel Burton–Miller-type integral equations for single-material and multi-material cases, respectively. After verifying the performance of the present formulations through various numerical experiments in Section 7, we conclude the paper and discuss future directions in Section 8.
@@ 2 Statement of the boundary value problem of interest
where $k_{i}:=\omega\sqrt{\varepsilon_{i}}$ is the wave number of the waves propagating in $\Omega_{i}$, and $\omega>0$ is the underlying angular frequency. Here, we assume the time dependency of all physical quantities in the time domain as $\Re[f(\boldsymbol{x})e^{-\mathrm{i}\omega t}]$, where $f$ is either $u$ or $w$, $\mathrm{i}$ is the imaginary unit, and $t$ is the time. The boundary conditions (2) and (3) on $\Gamma:=\bigcup_{i=2}^{M}\overline{\partial\Omega_{i}}$ define the jump relations in the total field and its flux. The flux here is defined as $w:=\frac{1}{\varepsilon_{i}}\nabla u\cdot\boldsymbol{n}$ with the material constants $\varepsilon_{i}$, where $\boldsymbol{n}$ is the unit normal on $\Gamma$. $\boldsymbol{n}$ is directed from $\Omega_{1}$ on $\partial\Omega_{1}$. The direction of $\boldsymbol{n}$ on the other boundaries is properly defined and fixed. This type of boundary value problem is often referred to as the transmission problem and is commonly found in engineering fields such as acoustics and electromagnetics. For example, a solution $u$ of the BVP (1)–(4) physically indicates the time-dependent out-of-plane component of a magnetic field, and $w$ corresponds to the in-plane tangential electric wave on the boundary. In this case, $\varepsilon_{i}$ represents the relative dielectric constant of a material filled in $\Omega_{i}$. Note that, although we keep ourselves to two-dimensional cases in this paper, our BEM formulation in the following sections is considered effective for 3D problems almost as is.
@@ 3 Design concepts for the present boundary integral formalism
Since the BVP (1)–(4) at hand is defined in an unbounded domain, the boundary integral method is a natural choice for solving it. There are various candidates for the integral formulations, and the choice may significantly affect the accuracy and efficiency of the practical computation. In this study, we ensure the accuracy by utilising the Burton–Miller method. In this section, we discuss the design concept of the integral formulation from the viewpoint of efficiency. Here, let us abstractly denote the integral equation as ${\cal A}y=b$ corresponding to the BVP, where $\cal A$ is an integral operator, $y$ and $b$ are respectively unknown and known density functions. The discretised version of the equation upon the collocation, i.e. the system of algebraic linear equations corresponding to the integral equation is denoted as $\mathsf{A}\boldsymbol{y}=\boldsymbol{b}$ with $\mathsf{A}$, $\boldsymbol{y}$, and $\boldsymbol{b}$ respectively being the coefficient matrix, and unknown and known vectors. Note that the discussions presented herein are specifically geared towards the collocation method. Substantial adjustments shall be required when applying the Galerkin BEM approach [24].
@@ 3.1 Integral equation of the second-kind
The efficiency of the BEM largely depends on the spectral property of $\cal A$. It is thus natural to try constructing the integral equation of the Fredholm second kind, i.e. exploring $\cal A$ such that $\cal A=\cal I+\cal K$ holds, where $\cal I$ and $\cal K$ are the identity and compact operators in an appropriate function space, respectively. The Müller formulation and the SBIE are categorised into this type. It may, unfortunately, be impossible to find such a well-conditioned $\cal A$ if one employs the Burton–Miller method.
@@ 3.2 Integral equation of the first-kind and its Calderon’s preconditioning
Another possibility may take the help of so-called preconditioning. Among the various preconditioning techniques, the Calderon preconditioning [14, 15] is recognised as one of the best for the BEM. This method converts (or preconditions) the original integral equation as ${\cal A}{\cal M}^{-1}z=b$ such that ${\cal A}{\cal M}^{-1}={\cal I}+{\cal K}$ holds, where $\cal M$ is the right preconditioning operator, and $z:={\cal M}y$ is an auxiliary function. Niino and Nishimura [16] demonstrated that, for transmission problems, one may construct, with the PMCHWT formulation, $\cal A$ and $\cal M$ such that ${\cal M}^{-1}\simeq{\cal A}\Leftrightarrow{\cal A}^{2}={\cal I}+{\cal K}$ holds. In the case that the collocation is used to discretise the integral operator ${\cal A}$, the resultant matrix $\mathsf{A}$ naturally succeeds the property; $\mathsf{A}^{2}$ approximately equals the identity matrix, and thus almost all eigenvalues of $\mathsf{A}^{2}$ exist only at a single point in the complex plane. They also pointed out that, once the integral operator is appropriately formulated in this manner, one does not have to explicitly multiply the (inverse of) preconditioner for the collocation BEM. This can be understood as follows: Let us consider the preconditioned system $\mathsf{A}^{2}\boldsymbol{z}=\boldsymbol{b}$ upon discretisation and solve this by the GMRES [9]. Let us first recall the fact that GMRES after $n$ iterations finds the best approximated solution of the algebraic equations from the following Krylov subspace:
@@
In summary, we just need to appropriately arrange the integral operator $\cal A$ such that it satisfies ${\cal A}^{2}=\cal I+\cal K$ and to solve the collocation-discretised linear equations without multiplying the preconditioner. The present paper is based on the discovery that we can construct such a $\cal A$ for the Burton–Miller-based integral equation even for transmission problems.
@@ 3.3 Spectral property of the conventional Burton–Miller method for transmission problems
where ${\cal I}$ is the identity operator. The integral operator in (14) is notoriously ill-conditioned. The upper and the lower diagonal parts are dominated by the hypersingular operator ${\cal N}$ and the weakly singular one ${\cal S}$, respectively. The eigenvalues of the former grow in magnitude, while those of the latter accumulate at the origin in the complex plane. This high contrast in the eigenvalues makes the resulting coefficient matrix considerably ill-conditioned. In Section 5, we modify the integral equation (14) with the help of the Calderon formulae to have a better conditioned one.
4 Calderon’s formulae and their collocation discretisation
Before deriving an well-conditioned Burton–Miller equations for transmission problems, let us review the properties of the integral operators (9)–(12). First, the operators ${\cal S}_{i}$, ${\cal D}_{i}$, and ${\cal N}_{i}$ are continuous mappings for $H^{-1/2}(\partial\Omega_{i})\rightarrow H^{1/2}(\partial\Omega_{i})$, $H^{1/2}(\partial\Omega_{i})\rightarrow H^{1/2}(\partial\Omega_{i})$, $H^{1/2}(\partial\Omega_{i})\rightarrow H^{-1/2}(\partial\Omega_{i})$, respectively, and ${\cal D}_{i}^{*}$ is the adjoint of ${\cal D}_{i}$ [30] for $i=1,\cdots,M$. If the boundary is smooth, the single- and double-layer operators are compact from $H^{1/2}(\partial\Omega_{i})$ into $H^{1/2}(\partial\Omega_{i})$ [30], and ${\cal N}_{i}$ is bounded. It is also well known that the operators have the following relations:
@@ 4 Calderon’s formulae and their collocation discretisation
$\frac{\cal I}{4}$ $=({\cal D}^{*}_{i})^{2}-{\cal N}_{i}{\cal S}_{i}.$ (18)
The relations (15)–(18) are known as the Calderon formulae. It is also known that the formulae are valid modulo a compact operator even when the wave numbers of two operators in a single term are different from each other [15].
Since the product of compact operators is also compact [30], from (15) and (18), one notices that ${\cal S}_{i}$ (resp. ${\cal N}_{i}$) is the pseudo inverse of ${\cal N}_{i}$ (resp. ${\cal S}_{i}$) up to the factor of $-1/4$ modulo a compact operator. This is only true for the case that $\partial\Omega_{i}$ is smooth. However, for the collocation-discretised matrices $\mathsf{S}_{i}$ and $\mathsf{N}_{i}$ with constant elements respectively from the operators ${\cal S}_{i}$ and ${\cal N}_{i}$, the matrix products $\mathsf{S}_{i}\mathsf{N}_{i}$ and $\mathsf{N}_{i}\mathsf{S}_{i}$ approximates the identity up to the constant factor even when the boundary includes some corner points. This is because, in the collocation discretisation, the collocation points never get on any corner point; the collocation point is usually set on the centre of a boundary element. To numerically elaborate this fact, we computed the boundary element matrices $\mathsf{S}$ and $\mathsf{N}$ obtained as the collocation discretised version of the above operators on a square of unit side length. In the computation, each side of the square is divided into 100 line segments of equal length, the collocation points are set on the centre of boundary elements, and the wave number is set as $k=1.5$. The magnitude of element-wise absolute error of $-\mathsf{SN}$ and $-\mathsf{NS}$ to the identity matrix scaled by $1/4$ are at most $5.394\times 10^{-2}$ and $6.213\times 10^{-2}$, respectively. Based on this observation, it is concluded that the collocation boundary element matrices naturally inherit the spectral properties of integral operators.
5 The present formulation: A single material case
@@
${\cal A}_{1}^{2}$ $=\begin{pmatrix}\frac{1}{4}{\cal I}+\alpha\varepsilon_{2}{\cal S}_{2}{\cal N}_{1}&0\\ \frac{1+\alpha}{4}{\cal I}+\frac{\alpha(1+\alpha)}{2}{\cal N}_{1}-\alpha({\cal N}_{1}{\cal D}_{2}+\alpha{\cal D}_{1}^{*}{\cal N}_{1})-\alpha{\cal S}_{1}{\cal N}_{1}&\frac{\alpha^{2}}{4}{\cal I}+\alpha\varepsilon_{2}{\cal N}_{1}{\cal S}_{2}\end{pmatrix}+{\cal K}$
@@
where we abusively used ${\cal K}$ to denote two compact operators different from each other. The last equality of (20) is due to the Calderon formulae (15)–(18). Although ${\cal A}_{1}^{2}$ still involves the hypersingular operator ${\cal N}_{1}$, owing to the block lower structure (except for a compact perturbation), its eigenvalue accumulates only at the following two points in the complex plane: $\frac{1-\alpha\varepsilon_{2}}{4}$ and $\frac{\alpha(\alpha-\varepsilon_{2})}{4}$, and they are away from the origin of the complex plane (Recall that the setting $\alpha=-\mathrm{i}/k_{1}$ is used here). With these observations, we may expect that the GMRES applied to the linear algebraic equations obtained by discretising (19) via collocation converges fast. Nonetheless, we cannot deny that the hypersingular operator ${\cal N}_{1}$ in ${\cal A}_{1}^{2}$ may affect the convergence of the linear solver.
To derive a better conditioned integral equation than (19), we multiply a complex constant to its first row as